Over a period of months, an adult male patient has taken thirteen blood tests for uric acid. The mean concentration was x = 5.35 mg/dl. The distribution of uric acid in healthy adult males can be assumed to be normal, with σ = 1.85 mg/dl.

(a) Find a 95% confidence interval for the population mean concentration of uric acid in this patient's blood. What is the margin of error? (Round your answers to two decimal places.)

lower limit    
upper limit    
margin of error   

(b) What conditions are necessary for your calculations?

(c) Find the sample size necessary for a 95% confidence level with maximal margin of error E = 1.14 for the mean concentration of uric acid in this patient's blood. (Round your answer up to the nearest whole number.)
Answer #1

Date: 1 solution). Given that, x = 5.35 6= 1185 n= 13 At as do confidence level, x=1-95 olo 2=1-0.95 =0.05 a = o.os - 00025 2E=1114 6=1.85 At as do confidence level the zis x=1- as olo L= 1-o.q5 =o.os = 009 = 0.025 Z2122 20.025 = 1.96 sample size = 1